# A Momentum Balance Correction to the Non-Conservative One-Fluid Formulation in Boiling Flows using Volume-of-Fluid

### v1

@@ Abstract
A proven methodology to solve multiphase flows is based on the one-fluid formulation of the governing equations, which treats the phase transition across the interface as a single fluid with varying properties and adds additional source terms to satisfy interface jump conditions, e.g., surface tension and mass transfer. Used interchangeably in the limit of non-evaporative flows, recent literature has formalized the inconsistencies that arise in the momentum balance of the non-conservative one-fluid formulation compared to its conservative counterpart when phase change is involved. This translates into an increased sensitivity of the numerical solution to the choice of formulation. Motivated by the fact that many legacy codes using the non-conservative one-fluid formulation have been extended to phase-change simulations, the inclusion of two corrective forces at the interface and a modification of the pressure-velocity solver with an additional predictor-projection step are shown to recover the exact momentum balance in the evaporative non-conservative one-fluid framework for low-viscosity incompressible flows. This has direct implications for obtaining a physically meaningful pressure field and is seen to affect the dynamics of two-phase flows. In the high-viscosity domain, the discretization of the viscous term introduces a momentum imbalance which is highly dependent on the chosen method to model the phase transition. Lastly, the need to develop sub-models for heat and mass transfer and for surface tension becomes evident since typical grid resolutions defined as “resolved” in the literature may not be enough to capture interfacial phenomena.
@@
volume-of-fluid , boiling flows , one-fluid formulation , momentum balance , phase change , non-conservative formulation , Fractional Step Method
@@ 1 Introduction
Conservative formulations of the one-fluid momentum equation, coupled to momentum-consistent advection schemes, successfully deal with flows that traditionally induced numerical difficulties (e.g., high density ratio flows) [10, 18, 17]. Yet, non-conservative formulations of the one-fluid momentum equation are commonly used in the literature. Mainly out of numerical convenience, they have been extended to incompressible two-phase flows undergoing phase change [8, 11, 12, 13, 14, 15]. Although mathematically equivalent to the conservative form in the continuous phase, the analytical derivation of the jump conditions for the non-conservative formulation differs from the exact jumps obtained with the conservative form when phase change occurs and, therefore, are physically inconsistent across the interface [14]. In other words, numerical integration of the non-conservative forms leads to an ill-defined momentum balance expressed in terms of an erroneous pressure field.
This inconsistency may seem negligible for practical purposes when convective speeds are large compared to the Stefan flow or when surface tension dominates over the momentum jump caused by the change of phase; thus, it is usually ignored. However, it may become important in, e.g., pressurized systems with fluids undergoing phase change close to the critical point as the surface tension coefficient and latent heat of vaporization drop [19]. These configurations may include droplet-laden or bubbly flows. Despite the correct velocity jump may be recovered, the momentum imbalance leads to a wrong pressure field, compromising the dynamical evolution of the flow and analyses based on pressure data, such as determining turbulence statistics (i.e., pressure fluctuations), studying the vorticity dynamics (i.e., calculating the baroclinic torque), or calculating forces around bodies (e.g., drag force on a droplet). These issues challenge the use of the one-fluid framework in such cases.
This work proposes a correction for the non-conservative form of the one-fluid momentum equation in terms of additional body forces in the context of the Continuum Surface Force (CSF) model [20] and a modified predictor-projection method to recover the correct momentum jump when phase change occurs and, consequently, obtain a physically meaningful pressure field. Moreover, this paper is also intended as a reflective exercise to highlight how the choice of numerical methods in two-phase solvers with phase change drastically affects the numerical solution.
@@ 2.1 Governing Equations
The governing equations for the motion and evaporation (or condensation) of a single-component fluid are the Navier-Stokes equations, given by the mass conservation or continuity equation (1), the momentum equation (2), and the energy transport equation (3). The variables $\rho$, $\mathbf{u}=(u,v,w)$, $p$, $\mu$, and $\mathbf{g}$ are, respectively, the density, velocity, pressure, dynamic viscosity and gravity, while $h$, $k$, and $T$ are the specific enthalpy, thermal conductivity and temperature. A Newtonian fluid is assumed under Stokes hypothesis; thus, the viscous stress tensor is $\mathbf{T}=\mu\big{[}\nabla\mathbf{u}+(\nabla\mathbf{u})^{T}-\tfrac{2}{3}(\nabla\cdot\mathbf{u})\mathbf{I}\big{]}$, where $\mathbf{I}$ is the identity tensor.
@@
$\dot{m}^{\prime}=\rho_{L}(\mathbf{u}_{L}-\mathbf{u}_{\Gamma})\cdot\mathbf{n}_{\Gamma}=\rho_{G}(\mathbf{u}_{G}-\mathbf{u}_{\Gamma})\cdot\mathbf{n}_{\Gamma}$ (4)
$[\rho_{G}\mathbf{u}_{G}(\mathbf{u}_{G}-\mathbf{u}_{\Gamma})-\rho_{L}\mathbf{u}_{L}(\mathbf{u}_{L}-\mathbf{u}_{\Gamma})]\cdot\mathbf{n}_{\Gamma}=-(p_{G}-p_{L})\mathbf{n}_{\Gamma}+[\mathbf{T}_{G}-\mathbf{T}_{L}]\cdot\mathbf{n}_{\Gamma}-\sigma\kappa\mathbf{n}_{\Gamma}$ (5)
@@
where $\dot{m}^{\prime}$ is the mass flux per unit area across the interface ($\dot{m}^{\prime}>0$ for vaporization), $\mathbf{u}_{\Gamma}=\mathbf{u}_{G}-\tfrac{\dot{m}^{\prime}}{\rho_{G}}\mathbf{n}_{\Gamma}=\mathbf{u}_{L}-\tfrac{\dot{m}^{\prime}}{\rho_{L}}\mathbf{n}_{\Gamma}$ the interface velocity, $\sigma$ the surface tension coefficient (assumed constant) and $\kappa=\nabla\cdot\mathbf{n}_{\Gamma}$ the interface curvature. Note the subscripts $L$ and $G$ indicate liquid and gas values at the interface. Further, $L$ and $G$ are also used to refer to the fluid properties in each single-phase region since they are assumed constant for the purpose of this work. Tangent to the interface, the velocity and shear stress are continuous. Rearranging Eq. (4), the velocity jump is given by $(\mathbf{u}_{G}-\mathbf{u}_{L})\cdot\mathbf{n}_{\Gamma}=\dot{m}^{\prime}(\rho_{G}^{-1}-\rho_{L}^{-1})$. Then, substitution of Eq. (4) into Eqs. (5) and (6) results in an equation for the pressure jump across the interface and an equation for the mass flux, given by
$p_{L}-p_{G}=\sigma\kappa+([\mathbf{T}_{L}-\mathbf{T}_{G}]\cdot\mathbf{n}_{\Gamma})\cdot\mathbf{n}_{\Gamma}+(\dot{m}^{\prime})^{2}(\rho_{G}^{-1}-\rho_{L}^{-1})$ (7)
$\dot{m}^{\prime}=\frac{[k_{G}\nabla T_{G}-k_{L}\nabla T_{L}]\cdot\mathbf{n}_{\Gamma}}{h_{LV}}$ (8)
@@ 2.2 Interface Capture with Volume-of-Fluid
The indicator function $X(\mathbf{x},t)$ is transported with the material interface (i.e., $DX/Dt=0$), which in the case of $\dot{m}^{\prime}=0$ is advected with the fluid. The discontinuous velocity when $\dot{m}^{\prime}\neq 0$ requires some rethinking of the advection approach to obtain consistent interface displacements. Some works advect the interface directly with $\mathbf{u}_{\Gamma}$ [10, 11, 35]. In contrast, the transport equation for $X(\mathbf{x},t)$ is defined here by Eq. (9) using a liquid velocity $\mathbf{u}_{L}$ and a source term to account for the liquid volume subtracted (or added) due to the vaporizing (condensing) interface. This term is described by a mass flux per unit volume $\dot{m}=\dot{m}^{\prime}\delta_{\Gamma}$, active only at interface cells [13, 16, 17]. $\mathbf{u}_{L}$ is a divergence-free velocity field representative of the liquid phase obtained by subtracting the Stefan flow from $\mathbf{u}$ [11]. The use of $\mathbf{u}_{L}$ and $\rho_{L}$ in Eq. (9) is applicable to droplet-laden flows, whereas the subtraction of the Stefan flow results in a divergence-free gas phase velocity $\mathbf{u}_{G}$ in bubbly flows [11]. Consequently, $\mathbf{u}_{G}$ and $\rho_{G}$ replace $\mathbf{u}_{L}$ and $\rho_{L}$ in Eq. (9).
$\frac{\partial X}{\partial t}+\nabla\cdot(X\mathbf{u}_{L})-X\nabla\cdot\mathbf{u}_{L}=-\frac{\dot{m}}{\rho_{L}}$ (9)
@@ 2.3 Non-Conservative One-Fluid Formulation of the Navier-Stokes Equations
$\nabla\cdot\mathbf{u}=\dot{m}(\rho_{G}^{-1}-\rho_{L}^{-1})=\dot{m}^{\prime}(\rho_{G}^{-1}-\rho_{L}^{-1})\delta_{\Gamma}$ (10)
@@
The momentum jump across the interface resulting from Eq. (11) is given by [14]
$\rho(\mathbf{u}_{L}-\mathbf{u}_{G})\mathbf{u}_{\Gamma}\cdot\mathbf{n}_{\Gamma}+\rho\mathbf{u}\cdot(\mathbf{u}_{G}-\mathbf{u}_{L})\mathbf{n}_{\Gamma}=-(p_{G}-p_{L})\mathbf{n}_{\Gamma}+[\mathbf{T}_{G}-\mathbf{T}_{L}]\cdot\mathbf{n}_{\Gamma}-\sigma\kappa\mathbf{n}_{\Gamma}$ (12)
@@
where $\rho$ and $\mathbf{u}$ are the one-fluid density and velocity, respectively. The terms on the left-hand side (LHS) of Eq. (12) differ from Eq. (5), demonstrating the ill-defined momentum balance introduced by the numerical integration of Eq. (11) if $\dot{m}^{\prime}\neq 0$. Similarly, a different momentum balance is obtained if the non-linear term $\mathbf{u}\cdot\nabla\mathbf{u}$ is replaced by $\nabla\cdot(\mathbf{uu})-\mathbf{u}\nabla\cdot\mathbf{u}$. Note that the analysis presented in [14] is analytical in nature and does not include issues that may arise from a given discretization approach.
@@
This imbalance is addressed by aiming at the replacement of the LHS of Eq. (12) with that of Eq. (5) in the non-conservative formulation. The first term on the LHS of Eq. (12), related to the time derivative of Eq. (11) and the displacement of the interface, is taken care of during the predictor-projection method used in the flow solver (more details in Section 3.3). Then, two body forces similar to $\mathbf{f}_{\sigma}$ are added to the right-hand side (RHS) of Eq. (11): (1) $\mathbf{f_{\text{NC}}}$ to correct the momentum jump error introduced by the convective term in the non-conservative (NC) formulation, i.e., the second term on the LHS of Eq. (12); and (2) $\mathbf{f}_{\dot{m}^{\prime}}$ to impose the exact momentum jump induced by phase change, i.e., the LHS of Eq. (5), similar to [12]. These are given by
$\mathbf{f_{\text{NC}}}=\rho\dot{m}^{\prime}(\rho_{G}^{-1}-\rho_{L}^{-1})(\mathbf{u}\cdot\mathbf{n}_{\Gamma})\mathbf{n}_{\Gamma}\delta_{\Gamma}$ (13)
$\mathbf{f}_{\dot{m}^{\prime}}=-(\dot{m}^{\prime})^{2}(\rho_{G}^{-1}-\rho_{L}^{-1})\mathbf{n}_{\Gamma}\delta_{\Gamma}$ (14)
@@
where $\mathbf{f_{\text{NC}}}$ is obtained by rewriting $\rho\mathbf{u}\cdot(\mathbf{u}_{G}-\mathbf{u}_{L})\mathbf{n}_{\Gamma}$ using the relation $(\mathbf{u}_{G}-\mathbf{u}_{L})=\dot{m}^{\prime}(\rho_{G}^{-1}-\rho_{L}^{-1})\mathbf{n}_{\Gamma}$. Thus, Eq. (11) becomes
@@
$\rho\frac{\partial\mathbf{u}}{\partial t}+\rho(\mathbf{u}\cdot\nabla\mathbf{u})=-\nabla p+\nabla\cdot\bigg{(}\mu\big{[}\nabla\mathbf{u}+(\nabla\mathbf{u})^{T}\big{]}\bigg{)}+\rho\mathbf{g}+\mathbf{f}_{\sigma}+\mathbf{f}_{\dot{m}^{\prime}}+\mathbf{f_{\text{NC}}}$ (15)
@@
In this work, it is assumed that the pressure jump due to viscous stresses is small compared to the effects of surface tension or phase change. Thus, the approach relies on the fact that regularization of viscous forces across the interface is enough to recover the exact jump caused by the viscous term [37]. From a numerical standpoint, the jump introduced by the one-fluid velocity in the discretization of $\nabla\mathbf{u}$ in the viscous term, which scales with the inverse of the mesh size, is also deemed negligible. Nonetheless, this assumption merits further analysis and the proposed methods could be extended to include viscosity effects explicitly in the momentum jump, if necessary. Note that this numerical issue is independent of whether the conservative or non-conservative form of the momentum equation is used in the one-fluid framework with phase change. The inconsistent discretization of the viscous term dominates the discussion in Section 5.3.
@@
The aforementioned inconsistencies arising from the one-fluid formulation in the non-conservative form could be mitigated by using phase-wise velocities (i.e., $\mathbf{u}_{l}$ and $\mathbf{u}_{g}$) and recovering the exact momentum jump by means of body forces, e.g., Eq. (14), or by implementing a Ghost Fluid Method (GFM), e.g., [38]. However, precise control of the numerically induced jump may be lost with the first approach and consistent extrapolation of variables across the interface or the implementation of GFM are prone to increased numerical instabilities due to the non-smooth $\mathbf{n}_{\Gamma}$ distributions in the context of VOF. Thus, the one-fluid formulation with corrective terms may be preferable while seeking a final outcome similar to GFM. In contrast, GFM may be more suited for more diffuse methods such as Level-Set or Phase-Field.
@@
Possible inconsistencies must be carefully addressed in conservative formulations as well. Despite the analytical derivation of the jump conditions result in the exact momentum balance given by Eq. (5), the numerical approach can introduce issues. For example, during temporal integration or in the implementation of momentum-consistent advection schemes, such as that from [17], when dealing with a sharp velocity jump across the interface.

### v2

@@ Abstract
A proven methodology to solve multiphase flows is based on the one-fluid formulation of the governing equations, which treats the phase transition across the interface as a single fluid with varying properties and adds additional source terms to satisfy interface jump conditions, e.g., surface tension and mass transfer. Used interchangeably in the limit of non-evaporative flows, recent literature has formalized the inconsistencies that arise in the momentum balance of the non-conservative one-fluid formulation compared to its conservative counterpart when phase change is involved. This translates into an increased sensitivity of the numerical solution to the choice of formulation. Motivated by the fact that many legacy codes using the non-conservative one-fluid formulation have been extended to phase-change simulations, the inclusion of two corrective forces at the interface and a modification of the pressure-velocity solver with an additional predictor-projection step are shown to recover the exact momentum balance in the evaporative non-conservative one-fluid framework for low-viscosity incompressible flows. This has direct implications for obtaining a physically meaningful pressure jump across the interface and is seen to affect the dynamics of two-phase flows. In the high-viscosity domain, the discretization of the viscous term introduces a momentum imbalance which is highly dependent on the chosen method to model the phase transition. In the context of film boiling, this imbalance affects the time scales for the instability growth. Lastly, the need to develop sub-models for heat and mass transfer and for surface tension becomes evident since typical grid resolutions defined as “resolved” in the literature may not be enough to capture interfacial phenomena.
@@
volume-of-fluid , boiling flows , one-fluid formulation , momentum balance , phase change , non-conservative formulation
journal: Journal of Computational Physics
@@ 1 Introduction
Conservative formulations of the one-fluid momentum equation, coupled to momentum-consistent advection schemes, successfully deal with flows that traditionally induced numerical difficulties (e.g., high density ratio flows) [10, 18, 17]. Yet, non-conservative formulations of the one-fluid momentum equation are commonly used in the literature. Mainly out of numerical convenience, they have been extended to incompressible two-phase flows undergoing phase change [8, 11, 12, 13, 14, 15]. Although mathematically equivalent to the conservative form in the continuous phase, the analytical derivation of the jump conditions for the non-conservative formulation differs from the exact jumps obtained with the conservative form when phase change occurs and, therefore, are physically inconsistent across the interface [14]. In other words, numerical integration of the non-conservative forms leads to an ill-defined momentum balance expressed in terms of a non-physical pressure field.
This inconsistency may seem negligible for practical purposes when convective speeds are large compared to the Stefan flow or when surface tension dominates over the momentum jump caused by the change of phase; thus, it is usually ignored. However, it may become important in, e.g., pressurized systems with fluids undergoing phase change close to the critical point as the surface tension coefficient and latent heat of vaporization drop [19]. These configurations may include droplet-laden or bubbly flows. Despite the correct velocity jump may be recovered, the momentum imbalance leads to a pressure solution that compromises the dynamical evolution of the flow and analyses based on pressure data, such as determining turbulence statistics (i.e., pressure fluctuations), studying the vorticity dynamics (i.e., calculating the baroclinic torque), or calculating forces around bodies (e.g., drag force on a droplet). These issues challenge the use of the one-fluid framework in such cases.
This work proposes a correction for the non-conservative form of the one-fluid momentum equation in terms of additional body forces in the context of the Continuum Surface Force (CSF) model [20] and a modified predictor-projection method to recover the correct momentum balance when phase change occurs and, consequently, obtain a physically meaningful pressure jump. Moreover, this paper is also intended as a reflective exercise to highlight how the choice of numerical methods in two-phase solvers with phase change drastically affects the numerical solution.
@@ 2.1 Governing Equations
The governing equations for the motion and evaporation (or condensation) of a single-component fluid are the Navier-Stokes equations, given by the mass conservation or continuity equation (1), the momentum equation (2), and the energy transport equation (3). The variables $\rho$, $\mathbf{u}=(u,v,w)$, $p$, $\mu$, and $\mathbf{g}$ are, respectively, the density, velocity, pressure, dynamic viscosity and gravity, while $h$, $k$, and $T$ are the specific enthalpy, thermal conductivity and temperature. A Newtonian fluid is assumed under Stokes hypothesis; thus, the viscous stress tensor is $\mathbf{\tau}=\mu\big{[}\nabla\mathbf{u}+(\nabla\mathbf{u})^{T}-\tfrac{2}{3}(\nabla\cdot\mathbf{u})\mathbf{I}\big{]}$, where $\mathbf{I}$ is the identity tensor.
@@
$\dot{m}^{\prime\prime}=\rho_{L}(\mathbf{u}_{L}-\mathbf{u}_{\Gamma})\cdot\mathbf{n}_{\Gamma}=\rho_{G}(\mathbf{u}_{G}-\mathbf{u}_{\Gamma})\cdot\mathbf{n}_{\Gamma}$ (4)
$[\rho_{G}\mathbf{u}_{G}(\mathbf{u}_{G}-\mathbf{u}_{\Gamma})-\rho_{L}\mathbf{u}_{L}(\mathbf{u}_{L}-\mathbf{u}_{\Gamma})]\cdot\mathbf{n}_{\Gamma}=-(p_{G}-p_{L})\mathbf{n}_{\Gamma}+[\mathbf{\tau}_{G}-\mathbf{\tau}_{L}]\cdot\mathbf{n}_{\Gamma}-\sigma\kappa\mathbf{n}_{\Gamma}$ (5)
@@
where $\dot{m}^{\prime\prime}$ is the mass flux per unit area across the interface ($\dot{m}^{\prime\prime}>0$ for vaporization), $\mathbf{u}_{\Gamma}=\mathbf{u}_{G}-\tfrac{\dot{m}^{\prime\prime}}{\rho_{G}}\mathbf{n}_{\Gamma}=\mathbf{u}_{L}-\tfrac{\dot{m}^{\prime\prime}}{\rho_{L}}\mathbf{n}_{\Gamma}$ the interface velocity, $\sigma$ the surface tension coefficient (assumed constant) and $\kappa=\nabla\cdot\mathbf{n}_{\Gamma}$ the interface curvature. Note the subscripts $L$ and $G$ indicate liquid and gas values at the interface. Further, $L$ and $G$ are also used to refer to the fluid properties in each single-phase region since they are assumed constant for the purpose of this work. Tangent to the interface, the velocity and shear stress are continuous. Rearranging Eq. (4), the velocity jump is given by $(\mathbf{u}_{G}-\mathbf{u}_{L})\cdot\mathbf{n}_{\Gamma}=\dot{m}^{\prime\prime}(\rho_{G}^{-1}-\rho_{L}^{-1})$. Then, substitution of Eq. (4) into Eqs. (5) and (6) results in an equation for the pressure jump across the interface and an equation for the mass flux, given by
$p_{L}-p_{G}=\sigma\kappa+([\mathbf{\tau}_{L}-\mathbf{\tau}_{G}]\cdot\mathbf{n}_{\Gamma})\cdot\mathbf{n}_{\Gamma}+(\dot{m}^{\prime\prime})^{2}(\rho_{G}^{-1}-\rho_{L}^{-1})$ (7)
$\dot{m}^{\prime\prime}=\frac{[k_{G}\nabla T_{G}-k_{L}\nabla T_{L}]\cdot\mathbf{n}_{\Gamma}}{h_{LV}}$ (8)
@@ 2.2 Interface Capture with Volume-of-Fluid
The indicator function $X(\mathbf{x},t)$ is transported with the material interface (i.e., $DX/Dt=0$), which in the case of $\dot{m}^{\prime\prime}=0$ is advected with the fluid. The discontinuous velocity when $\dot{m}^{\prime\prime}\neq 0$ requires some rethinking of the advection approach to obtain consistent interface displacements. Some works advect the interface directly with $\mathbf{u}_{\Gamma}$ [10, 11, 35]. In contrast, the transport equation for $X(\mathbf{x},t)$ is defined here by Eq. (9) using a liquid velocity $\mathbf{u}_{L}$ and a source term to account for the liquid volume subtracted (or added) due to the vaporizing (condensing) interface. This term is described by a mass flux per unit volume $\dot{m}^{\prime\prime\prime}=\dot{m}^{\prime\prime}\delta_{\Gamma}$, active only at interface cells [13, 16, 17]. $\mathbf{u}_{L}$ is a divergence-free velocity field representative of the liquid phase obtained by subtracting the Stefan flow from $\mathbf{u}$ by solving a Poisson equation efficiently using an FFT-based solver [11]. Alternatively, one may use an iterative solver to solve the Poisson equation only in a narrow band around the interface [13]. The use of $\mathbf{u}_{L}$ and $\rho_{L}$ in Eq. (9) is applicable to droplet-laden flows, whereas the subtraction of the Stefan flow results in a divergence-free gas phase velocity $\mathbf{u}_{G}$ in bubbly flows [11]. Consequently, $\mathbf{u}_{G}$ and $\rho_{G}$ replace $\mathbf{u}_{L}$ and $\rho_{L}$ in Eq. (9).
$\frac{\partial X}{\partial t}+\nabla\cdot(X\mathbf{u}_{L})-X\nabla\cdot\mathbf{u}_{L}=-\frac{\dot{m}^{\prime\prime\prime}}{\rho_{L}}$ (9)
@@ 2.3 Non-Conservative One-Fluid Formulation of the Navier-Stokes Equations
$\nabla\cdot\mathbf{u}=\dot{m}^{\prime\prime\prime}(\rho_{G}^{-1}-\rho_{L}^{-1})=\dot{m}^{\prime\prime}(\rho_{G}^{-1}-\rho_{L}^{-1})\delta_{\Gamma}$ (10)
@@
A methodology to properly derive jump conditions in one-fluid formulations is given by [14]. The momentum jump across the interface resulting from Eq. (11) [14] is
$\rho(\mathbf{u}_{L}-\mathbf{u}_{G})\mathbf{u}_{\Gamma}\cdot\mathbf{n}_{\Gamma}+\rho\mathbf{u}\cdot(\mathbf{u}_{G}-\mathbf{u}_{L})\mathbf{n}_{\Gamma}=-(p_{G}-p_{L})\mathbf{n}_{\Gamma}+[\mathbf{\tau}_{G}-\mathbf{\tau}_{L}]\cdot\mathbf{n}_{\Gamma}-\sigma\kappa\mathbf{n}_{\Gamma}$ (12)
@@
where $\rho$ and $\mathbf{u}$ are the one-fluid density and velocity, respectively. The terms on the left-hand side (LHS) of Eq. (12) differ from Eq. (5), demonstrating the ill-defined momentum balance introduced by the numerical integration of Eq. (11) if $\dot{m}^{\prime\prime}\neq 0$. Note that the analysis presented in [14] is analytical in nature and does not include issues that may arise from a given discretization approach. Similarly, a different momentum balance is obtained if the non-linear term $\mathbf{u}\cdot\nabla\mathbf{u}$ is replaced with $\nabla\cdot(\mathbf{uu})-\mathbf{u}(\nabla\cdot\mathbf{u})$ ($\equiv\nabla\cdot(\mathbf{uu})$ in an incompressible flow), which also does not satisfy the form of Eq. (5). Still, the same steps outlined in the following lines to correct the momentum balance for the non-conservative form could be implemented to find the correction terms of this other form of the momentum equation.
@@
The imbalance is addressed by aiming at the replacement of the LHS of Eq. (12) with that of Eq. (5). The first term on the LHS of Eq. (12), arising from the time derivative of Eq. (11) due to the displacement of the interface, is taken care of during the predictor-projection method in the flow solver (see Section 3.3). Then, two body forces similar to $\mathbf{f}_{\sigma}$ are added to the right-hand side (RHS) of Eq. (11): (1) $\mathbf{f_{\text{NC}}}$ to cancel the momentum jump error introduced by the convective term in the non-conservative (NC) formulation, i.e., the second term on the LHS of Eq. (12); and (2) $\mathbf{f}_{\dot{m}^{\prime\prime}}$ to impose the exact momentum jump induced by phase change, i.e., the LHS of Eq. (5), similar to [12]. These are given by
$\mathbf{f_{\text{NC}}}=\rho\dot{m}^{\prime\prime}(\rho_{G}^{-1}-\rho_{L}^{-1})(\mathbf{u}\cdot\mathbf{n}_{\Gamma})\mathbf{n}_{\Gamma}\delta_{\Gamma}$ (13)
$\mathbf{f}_{\dot{m}^{\prime\prime}}=-(\dot{m}^{\prime\prime})^{2}(\rho_{G}^{-1}-\rho_{L}^{-1})\mathbf{n}_{\Gamma}\delta_{\Gamma}$ (14)
@@
where $\mathbf{f_{\text{NC}}}$ is obtained by rewriting $\rho\mathbf{u}\cdot(\mathbf{u}_{G}-\mathbf{u}_{L})\mathbf{n}_{\Gamma}$ using the relation $(\mathbf{u}_{G}-\mathbf{u}_{L})=\dot{m}^{\prime\prime}(\rho_{G}^{-1}-\rho_{L}^{-1})\mathbf{n}_{\Gamma}$. Thus, Eq. (11) becomes
@@
$\rho\frac{\partial\mathbf{u}}{\partial t}+\rho(\mathbf{u}\cdot\nabla\mathbf{u})=-\nabla p+\nabla\cdot\bigg{(}\mu\big{[}\nabla\mathbf{u}+(\nabla\mathbf{u})^{T}\big{]}\bigg{)}+\rho\mathbf{g}+\mathbf{f}_{\sigma}+\mathbf{f}_{\dot{m}^{\prime\prime}}+\mathbf{f_{\text{NC}}}$ (15)
If the convective term is written as $\nabla\cdot(\mathbf{uu})$, $\mathbf{f}_{\dot{m}^{\prime\prime}}$ is still identical but not $\mathbf{f_{\text{NC}}}$. For the term $\nabla\cdot(\mathbf{uu})$, the jump across $\Gamma$ is given by $\rho(\mathbf{u}_{G}\mathbf{u}_{G}-\mathbf{u}_{L}\mathbf{u}_{L})\cdot\mathbf{n}_{\Gamma}$ [14], which can be used to derive the necessary body force that cancels the momentum balance error at the interface, similar to $\mathbf{f_{\text{NC}}}$. A validation of whether the correction is achieved at a discrete level needs to be performed, which is show in Section 3.5 for the non-conservative form. Note that the imbalance caused by the temporal term across the interface is not corrected by means of a body force, i.e., it is not included in $\mathbf{f_{\text{NC}}}$. A correction at a discrete level using such approach is not straightforward given the tight coupling with the predictor-projection method.
@@
In this work, it is assumed that the pressure jump due to viscous stresses is small compared to the effects of surface tension or phase change. Thus, regularization of viscous forces is enough to recover the effects of the viscous term [37]. From a numerical standpoint, the jump introduced by the one-fluid velocity in the discretization of $\nabla\mathbf{u}$ in the viscous term, which scales with the inverse of the mesh size, is also deemed negligible. Nonetheless, this assumption merits further analysis and the proposed methods could be extended to include viscosity effects explicitly in the momentum jump, if necessary. This numerical issue is independent of the form of the momentum equation used in the one-fluid framework. The problematic discretization of the viscous term dominates the discussion in Section 6.
@@
The aforementioned inconsistencies could be mitigated by using phase-wise velocities (i.e., $\mathbf{u}_{l}$ and $\mathbf{u}_{g}$) and recovering the exact momentum jump by means of body forces, e.g., Eq. (14), or by implementing a Ghost Fluid Method (GFM) [38]. However, precise control of the numerically induced jump may be lost with the first approach and consistent extrapolation of variables across the interface or the implementation of GFM are prone to increased numerical instabilities due to the non-smooth $\mathbf{n}_{\Gamma}$ distributions in the context of VOF. Thus, the one-fluid formulation with corrective terms may be preferable while seeking a final outcome similar to GFM. In contrast, GFM may be more suited for diffuse methods such as Level-Set or Phase-Field.
@@
Possible inconsistencies must be carefully addressed in conservative formulations as well. Despite the analytical derivation of the jump conditions result in the exact momentum balance given by Eq. (5), the numerical approach can introduce issues. For example, during temporal integration or in the implementation of momentum-consistent advection schemes when dealing with a sharp velocity jump across the interface.